Solve the system of equations by the elimination method. (If the system is dependent, let y = c and enter a general solution in terms of c. If there is no solution, enter NO SOLUTION.)
5/6x-1/3y=-21
1/6x+2/3y=9
(x, y) =

Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

can clear fractions or deal with them directly
multiply the first by two
(10/6)x-(2/3)y=-42
(1/6)x+(2/3)y=9 add
(11/6)x=-33
multiply both sides by 6/11
x=-18
put in first equation and -15-1/3 y=-21
-1/3 y=-6
y=18
check in second
-3+12=9 checks
(-18, 18)

Answer by Alan3354(69443)

Where are we supposed to enter NO SOLUTION?